Given this information:
Expected demand during lead time = 295 units
Standard deviation of lead time demand = 20 units
Use Table.
Determine each of the following, assuming that lead time demand is distributed normally:

a. The ROP that will provide a risk of stockout of 2 percent during lead time. (Round your answer to the nearest whole number.)

ROP             units
  
b. The safety stock needed to attain a 2 percent risk of stockout during lead time. (Do not round intermediate calculations. Round your answer to the nearest whole number.)

Safety stock             units

c-1. Would a stockout risk of 1 percent require more or less safety stock than a 2 percent risk?

  • Less

  • More



c-2. Would the ROP be larger, smaller, or unaffected if the acceptable risk were 1 percent instead of 2 percent?
  

  • Smaller

  • Larger

  • Unaffected
Answer #1

A. ROP = (Demand during lead time) + (Z * standard deviation of lead time demand)

Z = 2.054 (for 2% stockout, using table)

Demand during lead time = 295

Standard deviation of lead time demand = 20

ROP = 295 + (2.054 * 20) = 336.08

B. Safety stock = Z * Standard deviation of lead time demand

= 2.054 * 20 = 41.08

C.1. Stockout risk of 1 percent require more safety stock than a 2 percent risk as value of Z increases as the risk decreases.

C.2. ROP will larger as Z will increase and therefore, ROP will increase.

Do give a positive rating if you think this answer helped